OpenStudy (anonymous):

Can i have a step by step for this. Differentiate the function. ((cos^2)t)

OpenStudy (richyw):

are you trying to differentiate \(\cos^2(t)\) ? you need to use the chain rule\[f(t)=(\cos t)^2\]so let \(u=\cos t\), then you have\[\frac{df}{dt}=\frac{df}{du}\frac{du}{dt}\]\[\frac{df}{dt}=\frac{d}{du}(u^2)\frac{d}{dt}(\cos t)\]\[=-2u\sin t\]\[=-2\cos t\sin t\]

OpenStudy (anonymous):

Yes thats wut im trying to diff.

OpenStudy (anonymous):

\[\cos ^{2}t \] is equal to (cos t)^2

OpenStudy (anonymous):

how do i knw when i hav to use the chain rule? why am i using the chain rule???